# Simulating Policy Impacts: Developing a Generative Scenario Writing Method to Evaluate the Perceived Effects of Regulation

### v1

@@ 5 Discussion and Further Work
In this work we developed a method utilizing LLMs to generate scenarios to evaluate policies that may mitigate negative impacts on society. We demonstrate this method by simulating scenarios both mitigated and non-mitigated by transparency legislation—Article 50 of the EU AI Act—to convey negative impacts of generative AI in the media ecosystem. We then asked human evaluators to rate these simulated futures in order to gauge human perception of the severity, plausibility, magnitude, and specificity to vulnerable populations of this policy mitigation. We believe this method offers an opportunity to conduct policy evaluation on a given set of impact types in a straightforward manner. We now recap our high level findings, discuss limitations and capabilities of this method, elaborate on immediate possibilities of future research, and situate this method in the context of risk-assessment work.
@@
Our study follows the goal of algorithmic impact assessment studies as it aims to map negative impacts of a given technology and explore mitigation strategies in how to prevent them (Selbst 2021; Moss et al. 2021; Guston 2014). Using scenario-writing as a bottom-up approach of illustrating detrimental impacts of AI technology (Amer, Daim, and Jetter 2013; Meßmer and Degeling 2023; Börjeson et al. 2006), we supplement the top-down driven landscape of AIA. In this study, we introduced LLMs as a cost-effective way to create noteworthy scenarios based on impact themes identified in a study with human respondents. To demonstrate this method, we evaluated 10 impact themes established in prior work by Kieslich, Diakopoulos, and Helberger (2023) as relevant to the impact of generative AI on the media and information ecosystem: autonomy, education, labor, legal rights, media quality, political, security, social cohesion, trustworthiness, and well-being. Plausibility scores for the scenarios we generated remained relatively high across the board, indicating that GPT-4 was able to produce scenarios of impacts that were perceived to be genuinely possible depictions of the near-future impact of generative AI in the information ecosystem. The introduced transparency legislation was perceived to mitigate depicted negative impacts (whether in terms of severity, magnitude, or specificity to vulnerable populations) with statistical significance for autonomy, labor, legal rights, media quality, political harms, trustworthiness, and well-being. It did not seem to be impactful in the areas of education, security, or social cohesion, suggesting that perhaps other policy interventions are needed to combat negative impacts in these areas. For one impact, media fatigue, data suggests that the policy could even increase the severity of the impact. There were only two areas in which the legislation was perceived as statistically significantly reducing the specificity of the harms for vulnerable populations: political and well-being negative impacts, although overall the impact types that we studied here were not perceived to be highly specific to vulnerable populations.
@@
This study demonstrates that LLMs can sufficiently simulate complex narratives about AI’s impact and mitigation strategies, further contributing to the ongoing exploration of the ability of LLMs to map underlying complex relationships of the world, adding to discourse amongst Park et al. (2023); Tu, Ma, and Zhang (2023); Kıcıman et al. (2023); Cai, Liu, and Song (2023). These results generally demonstrate that our approach for using LLMs to generate scenarios, use those scenarios to simulate new scenarios in light of a policy, and then evaluate perceptions of impacts using human raters, can be valuable for understanding the degree to which the policy may be effective and help understand trade-offs in how the policy relates to different impacts. Being able to rank the impacts that may be shifted as a result of a proposed policy could be used to then allocate scarce experimental resources towards verifying the potential of a policy to mitigate an impact, such as through policy pilot programs. Additionally, based on our findings, future work might devise concrete measures for specific, more severe, impacts and then measure them in the years before and after the EU AI Act comes into force.
@@
In this study we only explored one part of one piece of legislation, but other policy proposals such as US President Biden’s Executive Order on AI (Biden E.O. 14110 of Oct 30, 2023) or the Chinese Generative AI Regulation (of China No. 15, July 13, 2023) offer interesting potential points of comparison. Future work will need to extend this method to account for more complex policies as well as to consider how different policies may interact to produce impacts on different stakeholders in society (and in different societies). The scenarios we generated also largely represent a best-case situation for policy impact as they largely assume that the policy being evaluated would be effectively enforced and upheld, yet it is possible that only certain elements of the legislation would be adhered to or properly enforced.
@@
Our implementation of the method also had a specified timeframe: five years in the future. We do not know how plausible the scenarios would be at more immediate time intervals (e.g., 6 months to 1 year) or further out in the future. It’s possible the scenarios could take on more of a science fiction element if we allowed for longer time intervals, which could potentially diminish the plausibility of scenarios and thus their relevance to policy evaluation. Additionally, we did not specify the level of jurisdiction for which the legislation would be enacted and enforced—it could have differing effects if implemented at a local versus federal or global level or according to the specific policy environment or set of institutional actors in a jurisdiction, but we did not specify this in our generated scenarios.
We further position the method established in this work with respect to the US NIST Risk Management Framework (NIST 2023)—a structure that organizations can use to facilitate risk management of design, development, and deployment of AI tools. The framework organizes risk management into three steps: mapping, measuring, and management. Creating the typologies, such as the one utilized in this paper (Kieslich, Diakopoulos, and Helberger 2023) constitute the mapping step, and then this method introduces a way to both measure the (perception of) severity and magnitude of these impacts to allow for stakeholders in power (e.g., legislators, government officials, AI researchers) to manage these harms. This method can assist with the prioritization of managing these harms by exploring the effectiveness of potential mitigation strategies which assists in deciding which risks to address first. Measuring severity, plausibility, magnitude, and specificity to vulnerable populations can also inform mitigation strategies beyond policy intervention. Our method is straightforward to adapt and can be extended and tested in new domains.
@@
We want to further emphasize the limitations of this work that are derived from the use of GPT-4. Of the 50 impact types outlined by Kieslich, Diakopoulos, and Helberger (2023), we were only able to generate 39 utilizing our prompting method (Appendix A.2). When we framed the prompts more broadly in terms of the ten impact themes, the generated scenarios focused on a heavily skewed set of impact types (see Figure 5 in the Appendix A.3); for instance, well-being harms focused almost entirely on mental harm with little acknowledgement of addiction, physical harm, or reputation. Finally, we were entirely unable to generate some harms (e.g., discrimination), which we hypothesize could be due to the highly tuned nature of this model to prevent toxicity being present in outputs. There will be biases present in all LLMs—both due to tuning and training data. Future work needs to explore how those biases may be reflected in scenarios and in models’ simulation of how a policy might impact a scenario.
@@ A.3 Proportion of Time LLM Uncovered Human Labeled Impact Types
Figure 5: Heat map displaying how many times human-identified specific impact types from (Kieslich, Diakopoulos, and Helberger 2023) were present in 10 LLM generated scenarios for each general impact theme. We prompted GPT-4 to generate a scenario 10 times for each of the 10 impact themes: autonomy, education, labor, legal rights, media quality, political, security, social cohesion, trustworthiness, and well-being. We then had two of the authors independently analyze whether each of the specific impact types were present in the generated scenarios. Darker colors correspond to impact types that were present more times than those with lighter colors. White backgrounds correspond to impact types that were never present in the scenarios generated by prompts focusing on the general impact theme.

### v2

@@ 5 Discussion and Future Work
In this work we developed a method utilizing LLMs to generate scenarios to evaluate policies that may mitigate negative impacts on society. We demonstrate this method by simulating scenarios both mitigated and non-mitigated by transparency legislation (Article 50 of the EU AI Act) to convey negative impacts of generative AI in the media ecosystem. We then asked human evaluators to rate these simulated futures in order to gauge human perception of the severity, plausibility, magnitude, and specificity to vulnerable populations of this policy mitigation. This method offers an opportunity to conduct policy evaluation on a given set of impact types in a straightforward manner. We now recap our high level findings, discuss limitations and capabilities of this method, elaborate on immediate possibilities of future research, and situate this method in risk-assessment work.
@@
This study follows the goal of algorithmic impact assessment studies as it aims to map negative impacts of a given technology and explore mitigation strategies in how to prevent them (Selbst 2021; Moss et al. 2021). Using scenario-writing as a bottom-up approach of illustrating detrimental impacts of AI technology (Amer, Daim, and Jetter 2013; Meßmer and Degeling 2023), we supplement the top-down driven landscape of AIA. In this study, we introduced LLMs as a cost-effective way to create scenarios based on an already established impact typology (Kieslich, Diakopoulos, and Helberger 2024), evaluating scenarios across 10 impact themes relevant to the impact of generative AI on the media and information ecosystem: autonomy, education, labor, legal rights, media quality, political, security, social cohesion, trustworthiness, and well-being. Plausibility scores for the scenarios we generated remained relatively high across the board, indicating that GPT-4 was able to produce scenarios of impacts that were perceived to be genuinely possible depictions of the near-future impact of generative AI in the information ecosystem. The introduced transparency legislation was perceived to mitigate depicted negative impacts (whether in terms of severity, magnitude, or specificity to vulnerable populations) with statistical significance for autonomy, labor, legal rights, media quality, political harms, trustworthiness, and well-being. It did not seem to be impactful in the areas of education, security, or social cohesion, suggesting that perhaps other policy interventions are needed to combat negative impacts in these areas. For one specific impact, media fatigue, data suggests that the policy could even increase the severity of the impact.
@@
Our findings demonstrate that LLMs can sufficiently simulate narratives about AI’s impact and mitigation strategies, further contributing to the ongoing exploration of the ability of LLMs to map underlying complex relationships of the world, adding to discourse amongst Park et al. (2023); Tu, Ma, and Zhang (2023); Kıcıman et al. (2023); Cai, Liu, and Song (2023). By conducting a user study we are able to understand how individuals perceive the potential impacts of given policies—a proof-of-concept for a future more democratized process that would leverage a representative sample. Such input could inform a wider matrix of trade offs that policymakers take into account when drafting policy. These results generally demonstrate the efficacy of our approach using LLMs to generate scenarios, using those scenarios to simulate new scenarios in light of a policy, and then evaluating perceptions of impacts using human raters. We envision this process could be used by policy makers and researchers alike to elucidate some potential impacts and efficacy of a given piece of policy and to help understand the trade-offs in how the policy relates to different impacts. Before a policy is formally codified it could be used to workshop policies at a brainstorming stage and to stimulate potential avenues of how policies could turn out in practice. Being able to rank the impacts that may be impacted by a proposed policy could then be used to allocate scarce experimental resources towards verifying the potential of a policy to mitigate an impact, such as through policy pilot programs.
Crucial to future work will be evaluating the approach we prototype here with policymakers in order to assess whether there are gaps that need to be bridged to make it useful in practice. Prior work underlines the importance of involving non-experts and laypeople in anticipatory assessments (Bonaccorsi, Apreda, and Fantoni 2020; Metcalf et al. 2021). However, critical judgment and expertise are necessary to confidently incorporate that information into larger processes of policy development. This includes getting (policy) expert opinions that validate the plausibility of the described policy effects in light of complex interactions between law, technology, and society.
@@
In this study we only explored one part of one piece of legislation, but other policy proposals such as US President Biden’s Executive Order on AI (Biden E.O. 14110 of Oct 30, 2023) or the Chinese Generative AI Regulation (Cybersecurity Administration of China No. 15, July 13, 2023) offer interesting potential points of comparison. We acknowledge that we are exploring the potential impact of EU legislation in a US context—however the EU AI Act is a specific implementation of some general ideas around labeling and transparency of generative AI systems and those general ideas also exist in draft form in some proposed legislation in US Congress. We chose to use the EU AI Act as it was a specific expression of the general ideas which are similarly applicable in the US context. Future work will need to extend this method to account for more complex policies as well as to consider how different policies may interact to produce impacts on different stakeholders in society and in different societies.
@@
We further position the method established in this work with respect to the US NIST Risk Management Framework (NIST 2023)—a structure that organizations can use to facilitate risk management of design, development, and deployment of AI tools. The framework organizes risk management into three steps: mapping, measuring, and management. Creating the typologies, such as the one utilized in this paper constitute the mapping step, and then this method introduces a way to both measure the (perception of) severity and magnitude of these impacts to allow for stakeholders in power (e.g., legislators, government officials, AI researchers) to manage these harms. This method can assist with the prioritization of managing these harms by exploring the effectiveness of potential mitigation strategies which assists in deciding which risks to address first. Measuring severity, plausibility, magnitude, and specificity to vulnerable populations can also inform mitigation strategies beyond policy intervention. Our method is straightforward to adapt and can be extended and tested in new domains.
@@
We want to further emphasize the limitations of this work that are derived from the use of GPT-4. Of the 50 impact types outlined by the Kieslich taxonomy, we were only able to generate 39 utilizing our prompting method (Appendix A.2). When we framed the prompts more broadly in terms of the ten impact themes, the generated scenarios focused on a heavily skewed set of impact types (see Figure 5 in the Appendix A.3). Finally, we were entirely unable to generate some harms (e.g., discrimination), which we hypothesize could be due to the highly tuned nature of this model to prevent toxicity being present in outputs. There will be biases present in all LLMs—both due to tuning and training data. Future work needs to explore how those biases may be reflected in scenarios and in models’ simulation of how a policy might impact a scenario.
@@ 6 Conclusion
Acknowledgments
This research is partially supported by UL Research Institutes through the Center for Advancing Safety of Machine Intelligence.
@@ A.3 Proportion of Time LLM Uncovered Human Labeled Impact Types
Figure 5: Heat map displaying how many times human-identified specific impact types from (Kieslich, Diakopoulos, and Helberger 2024) were present in 10 LLM generated scenarios for each general impact theme. We prompted GPT-4 to generate a scenario 10 times for each of the 10 impact themes: autonomy, education, labor, legal rights, media quality, political, security, social cohesion, trustworthiness, and well-being. We then had two of the authors independently analyze whether each of the specific impact types were present in the generated scenarios. Darker colors correspond to impact types that were present more times than those with lighter colors. White backgrounds correspond to impact types that were never present in the scenarios generated by prompts focusing on the general impact theme.